OpenStudy (anonymous):

Enrollment at a local university is 8,500. There are one third as many students over the age of 25 as there are 25 and under. How many students are 25 and under and how many are over 25?

OpenStudy (anonymous):

6375 students are 25 and under 2125 students are over 25

OpenStudy (anonymous):

how did u get thet answer may u show me your work i do not understand

OpenStudy (anonymous):

Let x = students 25 and under (1/3)x + x = 8500 x + 3x = 25500 (Multiplied the equation above by 3 to clear the fraction) 4x = 25500 x = 6375 (1/3) of 6375 = 1/3 * 6375 = 2125

OpenStudy (anonymous):

thank u makes sense i see where i went wrong in my solving